### v2

# A fast and exact approach for stabilizer Rényi entropy via the XOR–FWHT algorithm

Abstract
Quantum advantage is widely understood to rely on key quantum resources beyond entanglement, among which nonstabilizerness (quantum “magic”) plays a central role in enabling universal quantum computation. However, a direct brute-force enumeration of all Pauli strings and the corresponding expectation values from a length-$2^{N}$ state vector, where $N$ is the system size, yields an overall computational cost scaling as $O(8^{N})$, which quickly becomes infeasible as the system size grows. Here we reformulate the second-order stabilizer Rényi entropy in a bitstring language, expose an underlying XOR-convolution structure on $\mathbb{Z}_{2}^{N}$, and reduce the computation to $2^{N}$ fast Walsh-Hadamard transforms of length, together with pointwise operations, yielding a deterministic and exact XOR fast Walsh-Hadamard transforms algorithm with runtime scaling $O(N4^{N})$ and natural parallelism. This algorithm enables high-precision, medium-scale exact calculations for generic state vectors. It provides a practical tool for probing the scaling, phase diagnostics, and dynamical fine structure of quantum magic in many-body systems.
I Introduction
Quantum advantage refers to the possibility that a quantum computer achieves an exponential speedup over classical computation for specific tasks. This capability is commonly understood within the framework of resource theories [RevModPhys.91.025001] in which a set of key quantum resources collectively underpins computational advantage [RevModPhys.91.025001, shor, NielsenChuang2010, Daley2022, huang2025vastworldquantumadvantage, lanes2025frameworkquantumadvantage]. Among the resources most closely tied to quantum advantage are entanglement [Entanglement1, Entanglement2] and nonstabilizerness [chitambar2019quantum, liu2022manybody, PhysRevLett.128.050402, Leone_2024, tirrito2024quantifying, Turkeshi_2023, turkeshi2025pauli, Turkeshi_2024_2, jasser2025stabilizerentropyentanglementcomplexity, viscardi2025interplayentanglementstructuresstabilizer, Iannotti2025entanglement, cusumano2025nonstabilizernessviolationschshinequalities, bittel2025operationalinterpretationstabilizerentropy, varikuti2025impactcliffordoperationsnonstabilizing, tirrito2025anticoncentrationnonstabilizernessspreadingergodic, zhang2025stabilizerrenyientropytransition, qian2025quantum, moca2025nonstabilizernessgenerationmultiparticlequantum, dowling2025magic, bera2025nonstabilizernesssachdevyekitaevmodel, masotllima2024stabilizer, aditya2025mpembaeffectsquantumcomplexity, hernándezyanes2025nonstabilizernessquantumenhancedmetrologicalprotocols, falcão2025magicdynamicsmanybodylocalized, sticlet2025nonstabilizernessopenxxzspin, tirrito2025universalspreadingnonstabilizernessquantum, zhang2024quantummagicdynamicsrandom, cao2025gravitationalbackreactionmagical, Tarabunga_2024, qiant25, qian2025, hang2025, qiant2024a, frau2025, fan2025disentangling, huang2024cliffordcircuitsaugmentedmatrix, ding2025evaluating, korbany2025longrangenonstabilizernessphasesmatter, wei2025longrangenonstabilizernessquantumcodes, tarabunga2025efficientmutualmagicmagic, szombathy2025independentstabilizerrenyientropy, hou2025stabilizerentanglementenhancesmagic, hoshino2025stabilizerrenyientropyconformal, tarabunga2024magictransitionmeasurementonlycircuits, tirrito2025magicphasetransitionsmonitored, wang2025magictransitionmonitoredfree, santra2025complexitytransitionschaoticquantum, Haug2025probingquantum, Haug2023stabilizerentropies, Haug_2023_1, Lami_2023_2, Lami_2024, tarabunga2023manybody, tarabunga2024nonstabilizerness], which is also known as quantum magic. These resources are widely regarded as necessary ingredients for universal quantum computation [PhysRevLett.110.060504], and they play a central role in the construction and optimization of practical quantum devices [PhysRevLett.100.030504, PhysRevLett.91.147902, PhysRevLett.93.040502, PRXQuantum.3.020328]. It is important to emphasize that strong entanglement alone does not guarantee quantum advantage, since stabilizer states can be highly entangled while remaining efficiently simulable on classical computers under the Gottesman-Knill theorem [gottesman1998heisenbergrepresentationquantumcomputers, GT]. Nonstabilizerness, therefore, serves as an independent resource that quantifies the departure from the stabilizer manifold and that reflects the consumption of non-Clifford resources required to prepare a given state, which is why it is viewed as a particularly scarce and crucial resource for universal quantum computation when non-Clifford operations are costly in experiments and in fault-tolerant architectures [Preskill2018NISQ, PRXQuantum.3.020333, Quantinuum2025, Aasen2025, Peham2025].
In recent years, the behavior of magic and related nonstabilizerness quantifiers in many-body quantum systems has also become an active topic at the interface of quantum information [HaydenPreskill2007JHEP, SekinoSusskind2008JHEP, HosurQiRobertsYoshida2016JHEP, vonKeyserlingkRakovszkyPollmannSondhi2018PRX, KhemaniVishwanathHuse2018PRX, Swingle2018NatPhys, LiuZhangYinZhang2024PRL, LiuZhangYinZhangYao2024arXiv, YuLiZhang2025CPL, HuangLiHuseChan2023Scholarpedia, PhysRevResearch.7.L012011, Chen2025subsystem, PhysRevB.110.064323, PhysRevLett.132.240402] and many-body physics [PhysRevA.106.042426, Tarabunga2024criticalbehaviorsof, PhysRevA.108.042407, PhysRevA.110.022436, Passarelli2025chaosmagicin, PhysRevB.110.045101, delafuente2025topologicalstabilizermodelscontinuous, PRXQuantum.4.040317, PhysRevB.111.L081102, 10.21468/SciPostPhys.18.5.165, PaviglianitiLamiColluraSilva2025PRXQ, PRXQuantum.5.030332, li2025nonstabilizernessstarkmanybodylocalization, xfp5-hhs4, 1jzy-sk9r, LarkinOvchinnikov1969JETP, ShenkerStanford2014JHEP, MaldacenaShenkerStanford2016JHEP, Swingle2018NatPhys, HashimotoMurataYoshii2017JHEP, RobertsYoshida2017JHEP, CotlerHunterJonesLiuYoshida2017JHEP, LeoneOlivieroZhouHamma2021Quantum, GotoNosakaNozaki2022PRD, PaviglianitiLamiColluraSilva2025PRXQ, Prosen2021Chaos, HuangLiHuseChan2023Scholarpedia, maity2025localspreadingstabilizerrenyi, aditya2025growthspreadingquantumresources, Li_2025, lhz3, lhz2, PhysRevB.111.184302, zhao2025entanglementphasesphasetransitions, PhysRevA.109.062226, li2025quantummpembaeffectlongranged, sxz1, sxz2, sxz3, sxz4, liu2025noisymonitoredquantumcircuits, 6722-tf9c, yu2025gaplesssymmetryprotectedtopologicalstates, Koh_2022, Koh_2024, Tianqi, shenruizhe, digital1, digital2, digital3]. Quantifying nonstabilizerness is challenging both conceptually and computationally. A broad range of magic measures has been proposed, which includes quantifiers based on quasiprobability representations, robustness-type measures, and measures related to operational or simulation overhead [Veitch2014, WOOTTERS19871, Veitch_2012, PhysRevLett116250501, Bravyi, PhysRevLett.124.090505, Heimendahl2021stabilizerextentis]. For generic many-body states, however, evaluating these quantifiers typically entails exponential computational cost, which severely limits their applicability to larger system sizes.
The stabilizer Rényi entropies (SREs) [PhysRevLett.128.050402], provide a complementary perspective for quantifying nonstabilizer resources [PhysRevLett.128.050402, PhysRevA.107.022429]. Since SREs form a family of rigorously established magic monotones [PhysRevA.110.L040403], they avoid the need to construct a nonstabilizer decomposition explicitly, and because they are expressed directly in terms of Pauli statistics, they provide an attractive bridge between theory, numerics, and measurement-based evaluation protocols. Among the different orders, the second-order SRE (2-SRE) is distinguished as the simplest nontrivial SRE and because it has an explicit polynomial structure expressible as a quartic sum of Pauli expectation values. For an $N$-qubit pure state $|\psi\rangle$ in a Hilbert space of dimension $d$, the 2-SRE is defined as
$\mathcal{M}_{2}(|\psi\rangle)=-\log\!\left(\frac{1}{d}\ \sum_{P\in\mathcal{P}_{N}}\bigl|\langle\psi|P|\psi\rangle\bigr|^{4}\right),$ (1)
where $\mathcal{P}_{N}$ denotes the set of $N$-qubit Pauli strings, which is understood modulo an overall phase since such phases do not affect the modulus in (1). This expression can be interpreted as a fourth-moment concentration of Pauli expectation values, so that $\mathcal{M}_{2}$ becomes smaller when the Pauli statistics are highly structured, as in stabilizer states, and becomes larger when the distribution is more spread, as in generic nonstabilizer states.
Despite the explicit form of (1), its direct numerical evaluation remains prohibitively expensive. The summation runs over an exponentially large set of Pauli strings whose cardinality scales as $|\mathcal{P}_{N}|\!\sim\!4^{N}$, and for a generic state-vector input, computing a single expectation value $\langle\psi|P|\psi\rangle$ typically requires a full summation over $d$ amplitudes. As a consequence, the brute-force enumerated approach (Hereinafter referred to as brute-force approach) is often characterized by an overall scaling of $O(8^{N})$, which in practice confines exact calculations for spin-$\frac{1}{2}$ systems to relatively small system sizes, usually $N\!\lesssim\!14$.This limitation is structural rather than technical and constitutes a central computational bottleneck in many-body studies of nonstabilizerness, as it severely restricts controlled access to scaling behaviour, phase diagnostics, and dynamical phenomena at larger $N$.
To overcome this bottleneck, several computational strategies have been developed for the 2-SRE and related quantities. (i). Tensor-network approaches can be practical when the state admits a low-entanglement representation or an effectively controlled bond dimension [Haug2023stabilizerentropies, PhysRevB.107.035148, PhysRevLett.131.180401, PhysRevLett.133.010601, PhysRevLett.133.010602, tarabunga2025efficientmutualmagicmagic, PRXQuantum.5.030313, liu2025stabilizerrenyientropytranslationinvariant, PRXQuantum.6.010345, hoshino2025stabilizerrenyientropyconformal, hoshino2025stabilizerrenyientropyencodes, PhysRevB.109.174207], and their accuracy is then limited primarily by the fidelity of the underlying state representation when truncations are required. (ii). Sampling-based and Monte-Carlo estimators [PhysRevB.111.085144, PRXQuantum.4.040317, Tarabunga2024magicingeneralized, crew2025learningmagicschwingermodel, sinibaldi2025nonstabilizernessneuralquantumstates] avoid the complete $4^{N}$ enumeration by trading deterministic accuracy for statistical uncertainty, and their efficiency is governed by variance control and autocorrelation properties. (iii). For states with special algebraic structure, such as Gaussian states [PhysRevA.108.042407, collura2025nonstabilizernessfermionicgaussianstates, mx8t-l4hf, s1, PhysRevA.106.042426, rajabpour2025stabilizershannonrenyiequivalenceexact, wang2025magictransitionmonitoredfree], one can exploit reduced descriptions in terms of covariance matrices to obtain substantially faster evaluations. However, the applicability is then restricted to those structured families. These methods are powerful within their respective regimes. Still, their applicability and accuracy degrade outside their domain, as they typically rely on low-entanglement structures, controlled-variance stochastic estimators, or special state manifolds. There is an urgent need for a 2-SRE numerical method that does not assume special structure, yet remains broadly applicable, deterministically accurate, and computationally efficient by exploiting the algebraic structure intrinsic to (1).
In this work, we address this gap by presenting a deterministic and exact algorithm for evaluating the 2-SRE of a generic $N$-qubit pure state given as a state vector. We systematically rewrite the Pauli-string summation in (1) in a bitstring form and expose a convolution and correlation structure over the group $\mathbb{Z}_{2}^{N}$, i.e., XOR-convolutions [Ueno_Ito_Todo_Inoue_Minematsu_Ishikawa_Homma_2025] and correlations over bitstrings, which allows us to leverage orthogonality identities and the convolution theorem associated with the Walsh-Hadamard transform (WHT). This yields an exact formula amenable to efficient numerical implementation, in which the computation reduces to repeated fast Walsh-Hadamard transforms (FWHTs), pointwise operations, and index reshuffling. Accordingly, we refer to the algorithm proposed in this work as the XOR-FWHT algorithm. The algorithm requires Hilbert space of dimension $d\!=\!2^{N}$ FWHTs of length $d$, achieves a runtime scaling of $O(N4^{N})$ with modest additional memory, and is naturally parallelizable over the shift index. This reduction yields a substantial improvement over brute-force Pauli enumeration and enables exact evaluation of the 2-SRE up to $N\!\sim\!20$ for generic state vectors. While fast Walsh-Hadamard methods and $\mathbb{Z}_{2}^{N}$ convolution structures have been widely used in algorithm design and related computations involving quantum operators [Georges_2025], we provide a dedicated, systematic derivation of the 2-SRE and an open-source implementation of the resulting method.
The structure of this work is organized as follows. In Sec. II, starting from Eq. (1), we present a detailed derivation. After representing Pauli strings as bitstrings, we systematically simplify the expression by means of XOR convolution and the WHT, ultimately arriving at a compact form that is well suited for numerical implementation and exhibits significantly improved computational efficiency. In Sec. III, we provide the pseudocode of the proposed XOR-FWHT algorithm, and present the corresponding Python implementation details in the Appendix, together with benchmarks of the actual runtime performance on multicore CPU architectures. In Sec. IV, we present a series of numerical applications, which verify the consistency of our algorithm with both theoretical predictions and brute-force approaches; moreover, the high numerical precision of the method enables it to resolve subtle effects arising from small parameter variations in physical models. Finally, in Sec. V, we summarize the main results of this work and discuss potential directions for future developments of the XOR-FWHT algorithm.
II Algebraic Reformulation of the $2$-SRE
In this section, we derive Eq. (1), which constitutes the foundation of the numerical algorithm.
Bitstring representation of Pauli strings. The set of single-qubit Pauli operators is $\{\mathbf{1},X,Y,Z\}$. Noting that $Y=iXZ$, and that the overall phase of a Pauli string does not affect $|\langle\psi|P|\psi\rangle|^{4}$, we may, without loss of generality, ignore the global phase of Pauli strings in what follows and retain only their $X^{x}Z^{z}$ structure.
Any $N$-qubit Pauli string can be uniquely written as
$P=X_{1}^{x_{1}}\cdots X_{N}^{x_{N}}Z_{1}^{z_{1}}\cdots Z_{N}^{z_{N}}\equiv X^{x}Z^{z},$ (2)
where $x:=(x_{1},\dots,x_{N}),\;z:=(z_{1},\dots,z_{N})\in\mathbb{Z}_{2}^{N}$. Therefore, summing over all Pauli strings $P\in\mathcal{P}_{L}$ is equivalent to summing over all $x,z\in\mathbb{Z}_{2}^{N}$.
Expansion of the quantum state in the computational basis. We write the quantum state in the computational bitstring basis as
$|\psi\rangle=\sum_{t\in\mathbb{Z}_{2}^{N}}\phi_{t}\,|t\rangle,$ (3)
where $|t\rangle$ denotes the computational basis state labeled by the length-$N$ bitstring $t=(t_{1},\dots,t_{N})$.
Action of Pauli operators on the computational basis. The actions of the Pauli operators $X^{x}$ and $Z^{z}$ on the computational basis are given by
$X^{x}|t\rangle$ $=|t\oplus x\rangle,$ (4)
$Z^{z}|t\rangle$ $=(-1)^{z\cdot t}\,|t\rangle,$ (5)
where $\oplus$ denotes bitwise addition modulo $2$, and $z\cdot t\!=\!\sum_{j=1}^{N}z_{j}t_{j}\pmod{2}$ is the binary inner product. Consequently,
$P|\psi\rangle=\sum_{t}\phi_{t}\,(-1)^{z\cdot t}\,|t\oplus x\rangle.$ (6)
Pauli expectation values and XOR-correlation structure. It then follows that
$\langle\psi|P|\psi\rangle=\sum_{t}(-1)^{z\cdot t}\,\phi_{t\oplus x}^{*}\phi_{t}.$ (7)
This expression exhibits an XOR-correlation structure, namely a weighted sum over correlations between the wavefunction amplitudes $\phi_{t}$ and their bit-shifted counterparts $\phi_{t\oplus x}$. In the following derivation, we first fix $x$ and perform the summation over $z$.
Fourth-order modulus expansion and Walsh-Hadamard orthogonality. Define
$g_{x}(t):=\phi_{t\oplus x}^{*}\phi_{t},$ (8)
and introduce
$A(z):=\sum_{t\in\mathbb{Z}_{2}^{N}}(-1)^{z\cdot t}\,g_{x}(t),$ (9)
so that
$|A(z)|^{4}=A(z)A^{*}(z)A(z)A^{*}(z).$ (10)
Expanding this expression explicitly yields
$|A(z)|^{4}$ $=\sum_{t_{1},t_{2},t_{3},t_{4}}(-1)^{z\cdot(t_{1}\oplus t_{2}\oplus t_{3}\oplus t_{4})}$
$\qquad\qquad\times g_{x}(t_{1})\,g_{x}^{*}(t_{2})\,g_{x}(t_{3})\,g_{x}^{*}(t_{4}).$ (11)
Since the definition of the SRE involves a sum over all $z\in\mathbb{Z}_{2}^{N}$, we can exploit the orthogonality relation of the Walsh-Hadamard transform defined on the group $\mathbb{Z}_{2}^{N}$,
$\sum_{z\in\mathbb{Z}_{2}^{N}}(-1)^{z\cdot s}=\begin{cases}2^{N},&s=0,\\ 0,&s\neq 0,\end{cases}$ (12)
to explicitly carry out the summation over $z$, obtaining
$\sum_{z}|A(z)|^{4}$ $=2^{N}\sum_{\begin{subarray}{c}t_{1},t_{2},t_{3},t_{4}\\ t_{1}\oplus t_{2}\oplus t_{3}\oplus t_{4}=0\end{subarray}}g_{x}(t_{1})\,g_{x}^{*}(t_{2})\,g_{x}(t_{3})\,g_{x}^{*}(t_{4}).$ (13)
The orthogonality of the Walsh-Hadamard transform projects the originally unconstrained fourfold sum onto the subspace satisfying $t_{1}\oplus t_{2}\oplus t_{3}\oplus t_{4}=0$. This XOR constraint is the key to the subsequent efficient numerical implementation.
XOR-convolution structure and dimensionality reduction. The constraint $t_{1}\oplus t_{2}\oplus t_{3}\oplus t_{4}=0$ allows one to eliminate one variable, for instance, by writing $t_{4}=t_{1}\oplus t_{2}\oplus t_{3}$. Eq. (13) can thus be rewritten as
$\sum_{z}|A(z)|^{4}$ $=2^{N}\sum_{t_{1},t_{2},t_{3}}g_{x}(t_{1})\,g_{x}^{*}(t_{2})\,g_{x}(t_{3})\,g_{x}^{*}(t_{1}\oplus t_{2}\oplus t_{3}).$ (14)
The above triple sum has a characteristic XOR convolution structure (see Appendix A.3). To make this explicit, we introduce the intermediate function
$C_{x}(s):=\sum_{t\in\mathbb{Z}_{2}^{N}}g_{x}(t)\,g_{x}^{*}(t\oplus s),$ (15)
where $s\in\mathbb{Z}_{2}^{N}$.
Using Eq. (15) and appropriately relabeling the summation variables, Eq. (14) can be further simplified to
$\sum_{z}|A(z)|^{4}$ $=2^{N}\sum_{s\in\mathbb{Z}_{2}^{N}}\left|\sum_{t\in\mathbb{Z}_{2}^{N}}g_{x}(t)\,g_{x}^{*}(t\oplus s)\right|^{2}$
$=2^{N}\sum_{s\in\mathbb{Z}_{2}^{N}}|C_{x}(s)|^{2}.$ (16)
Exploiting properties of the WHT. We apply the WHT to $C_{x}(s)$, and define $G_{x}:=\text{WHT}(g_{x})$. Using property A.4, we obtain
$\text{WHT}(C_{x})=\sqrt{2^{N}}G_{x}\cdot G_{x}^{*}=\sqrt{2^{N}}|G_{x}|^{2}.$ (17)
Applying the Parseval identity of the WHT A.5 to Eq. (16) yields
$\sum_{s}|C_{x}(s)|^{2}=\sum_{k}|\text{WHT}(C_{x})(k)|^{2}.$ (18)
This further leads to
$\sum_{s}|C_{x}(s)|^{2}={2^{N}}\sum_{k}|G_{x}(k)|^{4}.$ (19)
Then we obtain (there is also a factor of $2^{N}$ in Eq. (16))
$\sum_{P\in\mathcal{P}_{L}}\bigl|\langle\psi|P|\psi\rangle\bigr|^{4}=4^{N}\sum_{x,k}{|\text{WHT}(\phi_{x\oplus k}^{*}\phi_{x})|^{4}}.$ (20)
We note that the FWHT provides an efficient numerical implementation of the WHT. According to Appendix A.6, Eq. (20) can be rewritten as
$\sum_{P\in\mathcal{P}_{L}}\bigl|\langle\psi|P|\psi\rangle\bigr|^{4}=\sum_{x,k}{|\text{FWHT}(\phi_{x\oplus k}^{*}\phi_{x})|^{4}}.$ (21)
Final result. Therefore, Eq. (1) can be written as
$\mathcal{M}_{2}(|\psi\rangle)=-\log\!\left(\frac{1}{d}\sum_{x,k}{|\text{FWHT}(\phi_{x\oplus k}^{*}\phi_{x})|^{4}}\right),$ (22)
Eq. (22) is the final expression. In the next section, we will show how to use it to design Algorithm 1.
III XOR-FWHT Algorithm for the $2$-SREs
By using Eq. (22), we design the Algorithm 1. It is worth emphasizing that the bitstrings $x$ and $k$ are represented by their integer encodings in the numerical implementation, which is straightforward for a computer to handle. The implementation requires performing $d$ FWHTs, each of length $d$, leading to an overall time complexity of $O(2^{N}\times N2^{N})=O(N4^{N})$. The additional memory overhead originates from storing the reordered conjugate vector $\overline{\phi[\bm{x}\oplus k]}\quad(x=0,\ldots,d-1)$; this overhead does not accumulate with $k$ and is therefore acceptable. Moreover, the computations corresponding to different values of $k$ are mutually independent, making the algorithm naturally well-suited to parallel implementation. We provide implementation details in Python in Appendix B.
We benchmarked the performance of the algorithm on a workstation equipped with two Intel Xeon Platinum 8352V processors (36 cores per CPU, 72 cores in total). In the benchmarks, the input states were randomly generated Haar states obtained by normalizing complex Gaussian random vectors, and we measured the wall-clock time for a single evaluation of the 2-SRE. For different system sizes $N$, the typical runtimes are as follows: approximately $0.88$ seconds for $N=14$; about $12$ seconds for $N=16$; about $300$ seconds for $N=18$; and about $2$ hours for $N=20$. These numerical results are consistent with the theoretical time complexity $O(N4^{N})$ of the algorithm.
Algorithm 1 XOR-FWHT Algorithm of the $\mathcal{M}_{2}$
Input: normalized state vector $\phi[0],\ldots,\phi[d-1]$in the computational basis , where $d=2^{N}$.
Output: $\mathcal{M}_{2}$.
[…]
3:for $k=0$ to d1d-1 do
4:  construct length-$d$ complex array $G$ by
$G[x]\leftarrow\overline{\psi[x\oplus k]}\,\psi[x]\qquad(x=0,\ldots,d-1)$
5:  $\hat{G}\leftarrow\mathrm{FWHT}(G)$
6:  $r\leftarrow r+\sum_{u=0}^{d-1}|\hat{G}[u]|^{4}$
7:end for
8:return $\mathcal{M}_{2}\leftarrow-\log_{2}\!\bigl(r/d\bigr)$
IV Applications in Physical Examples
In this section, we numerically present several representative physical examples by using algorithm 1 for evaluating 2-SRE.
IV.1 Example 1. validation on Haar random states
First, we randomly sample Haar-random pure states from the Hilbert space according to the Haar measure. As typical quantum states, Haar-random states exhibit volume-law entanglement near the Page limit at the subsystem level, which causes tensor-network-based approaches to compute $\mathcal{M}_{2}$ to encounter an exponentially growing bond-dimension bottleneck as the system size increases. In contrast, the XOR-FWHT algorithm proposed in this work does not rely on the bond dimension of a tensor-network representation, but instead acts directly on the state vector itself, and therefore enables efficient evaluation of the magic measure $\mathcal{M}_{2}$ for Haar-random states.
As shown in Fig. 1(a), the numerical results are in good agreement with the theoretical prediction [PhysRevB.111.054301], thereby validating the correctness and applicability of the algorithm for highly entangled states.
Figure 1: (a) The average 2-SRE $\mathcal{M}_{2}$ of Haar-random pure states as a function of the system size $N$. For each NN, we independently sample 1010010-100 Haar-random states and compute the average. The solid line shows the theoretically predicted behavior $\log_{2}(2^{N}+3)-2$, which is in good agreement with the numerical results, indicating that $\mathcal{M}_{2}$ of Haar states grows approximately linearly with $N$. (b) Time evolution of $\mathcal{M}_{2}$ in the dynamics of the XXZ model, comparing results obtained from the brute-force approach and from the XOR-FWHT algorithm proposed in this work. The two methods yield identical numerical results throughout the entire time evolution, verifying the correctness and stability of the XOR-FWHT algorithm in dynamical calculations.
IV.2 Example 2. Hamiltonian quench dynamics
Figure 2: Time evolution of the 2-SRE $\mathcal{M}_{2}$ for N=16N\!=\!16 spins evolved under the TFIM with a longitudinal field. Starting from $100$ random product states, $\mathcal{M}_{2}(t)$ rapidly grows and saturates at a value slightly below the Haar-random benchmark (dashed line) within the accessible time window.
Next, we consider a one-dimensional XXZ spin-$\frac{1}{2}$ chain with the Hamiltonian
$H_{\mathrm{XXZ}}=\sum_{i=0}^{N-1}J\left(\,X_{i}X_{i+1}+\,Y_{i}Y_{i+1}\right)+\Delta\,Z_{i}Z_{i+1},$ (23)
[…]
We also consider the transverse-field Ising model with an additional longitudinal field (TFIM with longitudinal field), whose Hamiltonian is given by
$H_{\mathrm{TFIM+LF}}=-J\sum_{i=0}^{N-1}Z_{i}Z_{i+1}-h_{x}\sum_{i=0}^{N-1}X_{i}-h_{z}\sum_{i=0}^{N-1}Z_{i},$ (24)
where we set $J=1$, $h_{x}=1.5$, and $h_{z}=1.5$, with the system size $N=16$.
The initial state is chosen as a tensor product of random single-qubit states,
[…]
where $\theta_{i}\in[0,\pi]$ and $\phi_{i}\in[0,2\pi)$ are independent random variables. Although these initial states are unentangled, they are generally non-stabilizer states and thus carry local non-stabilizer resources. Starting from $100$ different random initial states, we perform real-time evolution using the Krylov subspace method, and compute the SRE $\mathcal{M}_{2}$ during the evolution, followed by averaging over the results. The resulting dynamical behavior is shown in Fig. 2. We observe that, within the system size and time window considered, $\mathcal{M}_{2}(t)$ grows rapidly with time and gradually approaches saturation, while its value always remains slightly below the corresponding Haar-random-state benchmark. Notably, in Fig. 4(b) of Ref. [y9r6-dx7p], this subtle deviation from the Haar value is not clearly resolved. We conjecture that this may be related to the intrinsic statistical errors of numerical evaluation methods based on finite sampling, which make it difficult to resolve such small deviations in $\mathcal{M}_{2}$ at that level of precision. In contrast, the exact XOR-FWHT algorithm employed in this work enables high-precision evaluation of $\mathcal{M}_{2}$ without resorting to stochastic sampling, thereby allowing such subtle but physically meaningful deviations to be clearly revealed. This provides a new numerical tool for further investigating the fine structure of magic resources in dynamical processes.
IV.3 Example 3. random unitary circuit quench dynamics
[…]
As illustrated in Fig. 3b, we numerically simulated this setup using TensorCircuit-NG [Zhang_2023] for system sizes $N=10,12,14,16$. For all system sizes considered, the SRE dynamics saturates at long times to the corresponding Haar value.
V Conclusion and Outlook
In this work, we propose a fast and exact algorithm for evaluating medium-scale 2-SRE $\mathcal{M}_{2}$ with a computational time complexity of $O(N4^{N})$. Starting from the definition of $\mathcal{M}_{2}$, we derive a fully deterministic reformulation based on XOR convolution and the WHT, without introducing any approximation, stochastic sampling, or additional assumptions on the structure of the quantum state throughout the derivation. The resulting expression can be efficiently implemented using the FWHT, leading to a substantial reduction in computational cost compared to brute-force approaches. We have made the complete algorithmic details and the corresponding Python implementation publicly available. Through several numerical examples, we demonstrate that our method accurately reproduces the theoretical expectation value of $\mathcal{M}_{2}$ for Haar-random states, and that it is in full numerical agreement with brute-force results for the system sizes where such comparisons are possible. More importantly, the algorithm achieves high numerical precision, allowing one to resolve small but physically meaningful deviations in $\mathcal{M}_{2}$, thereby enabling the investigation of the fine structure of non-stabilizer resources in quantum states and dynamics.
Looking ahead, several promising directions remain for further exploration. One direct application of the present algorithm is the systematic study of the dynamical behavior of $\mathcal{M}_{2}$ in a variety of interacting quantum many-body models, as well as its interplay with other physical quantities such as entanglement and operator spreading. From a computational perspective, combining the current framework with tensor-network representations is particularly appealing, as this may further reduce the effective computational cost for large systems or weakly entangled states. More broadly, we expect that the algorithmic ideas developed in this work can serve as building blocks for the efficient evaluation of higher-order stabilizer Rényi entropies and related resource measures.
Acknowledgements.
J.-X. Zhong was supported by the National Natural Science Foundation of China (Grant Nos. 12374046 and 11874316), the Shanghai Science and Technology Innovation Action Plan (Grant No. 24LZ1400800), the National Basic Research Program of China (Grant No. 2015CB921103), and the Program for Changjiang Scholars and Innovative Research Teams in Universities (Grant No. IRT13093). H.-Z. Li is supported by a China Scholarship Council Scholarship.
Appendix A Introduction to the WHT
To facilitate the reader’s understanding of the logical structure of the derivations presented in the main text, we summarize in this appendix the properties and features of the WHT used in this work.
A.1 Definition
We first define the WHT of a function $f:\{0,1\}^{N}\to\mathbb{C}$ as
$\hat{f}(k)=\frac{1}{\sqrt{2^{N}}}\sum_{x\in\{0,1\}^{N}}(-1)^{k\cdot x}f(x),$ (26)
where $k\cdot x$ denotes the mod-$2$ inner product.
A.2 Property of Involution and Inverse Transform
The WHT satisfies
$\mathrm{WHT}\bigl(\mathrm{WHT}(f)\bigr)=f.$ (27)
Therefore, the inverse transform is given by
$\mathrm{IWHT}(f)=\mathrm{WHT}(f).$ (28)
A.3 Properties of XOR-Convolution Theorem
We define the XOR convolution as
$(f*_{\oplus}g)(x)=\sum_{y\in\mathbb{Z}_{2}^{N}}f(y)\,g(x\oplus y),$ (29)
then
$\mathrm{WHT}(f*_{\oplus}g)=\sqrt{2^{N}}\mathrm{WHT}(f)\cdot\mathrm{WHT}(g),$ (30)
where the symbol $\cdot$ on the right-hand side denotes pointwise multiplication of vectors.
A.4 Property I: Compatibility with Complex Conjugation
For any complex-valued vector $f$, one has
$\mathrm{WHT}(f^{*})(k)=\bigl(\mathrm{WHT}(f)(k)\bigr)^{*}.$ (31)
Consequently, $\mathrm{WHT}(f)\,\mathrm{WHT}(f^{*})=|\mathrm{WHT}(f)|^{2}$.
A.5 Property II: Parseval’s Theorem
The WHT preserves the $\ell_{2}$ norm:
$\sum_{x\in\mathbb{Z}_{2}^{N}}|f(x)|^{2}=\sum_{k\in\mathbb{Z}_{2}^{N}}\bigl|\mathrm{WHT}(f)(k)\bigr|^{2}.$ (32)
A.6 Numerical Algorithm of FWTH
The WHT admits an efficient numerical implementation known as the FWHT. Rather than defining a new transform, the FWHT computes the WHT by exploiting the recursive structure of the Hadamard matrix. In our implementation, the FWHT evaluates the unnormalized transform, which is related to the normalized WHT by
$\mathrm{FWHT}(f)=\sqrt{2^{N}}\,\mathrm{WHT}(f).$ (33)
Concretely, the FWHT proceeds via iterative butterfly updates of the form
$(a,b)\mapsto(a+b,a-b),$ (34)
applied recursively over $N$ levels for functions defined on $\{0,1\}^{N}$, resulting in an overall computational complexity of $O(N2^{N})$. Implementation details are provided in Appendix B.
Appendix B Python Implementation of Algorithm 1
1import numpy as np
2
3def FWHT_inplace(a):
4 n = a.shape[0]
5 h = 1
6 while h < n:
7 for i in range(0, n, 2*h):
8 u = a[i:i+h]
9 v = a[i+h:i+2*h]
10 u[:], v[:] = u+v, u-v
11 h *= 2
12
13def M_func(psi):
14 d = psi.shape[0]
15 x = np.arange(d, dtype=np.int64)
16 total = 0.0
17 for k in range(d):
18 A = np.conjugate(psi[x^k]) * psi
19 FWHT_inplace(A)
20 total += (np.sum(np.abs(A) ** 4))
21 return -np.log2(total/d)
It should be emphasized that, in the practical implementation of this code, we also employ JIT and parallelization techniques, enabling us to fully utilize multicore CPU resources.